# Spec2VolCAMU-Net: A Spectrogram-to-Volume Model for EEG-to-fMRI Reconstruction based on Multi-directional Time-Frequency Convolutional Attention Encoder and Vision-Mamba U-Net

### v1

@@ Abstract
High-resolution functional magnetic resonance imaging (fMRI) is essential for mapping human brain activity; however, it remains costly and logistically challenging. If comparable volumes could be generated directly from widely available scalp electroencephalography (EEG), advanced neuroimaging would become significantly more accessible. Existing EEG-to-fMRI generators rely on plain CNNs that fail to capture cross-channel time-frequency cues or on heavy transformer/GAN decoders that strain memory and stability. We propose Spec2VolCAMU-Net, a lightweight spectrogram-to-volume generator that confronts these issues via a Multi-directional Time-Frequency Convolutional Attention Encoder, stacking temporal, spectral and joint convolutions with self-attention, and a Vision-Mamba U-Net decoder whose linear-time state-space blocks enable efficient long-range spatial modelling. Trained end-to-end with a hybrid SSI-MSE loss, Spec2VolCAMU-Net achieves state-of-the-art fidelity on three public benchmarks, recording SSIMs of 0.693 on NODDI, 0.725 on Oddball and 0.788 on CN-EPFL, representing improvements of 14.5%, 14.9%, and 16.9% respectively over previous best SSIM scores. Furthermore, it achieves competitive PSNR scores, particularly excelling on the CN-EPFL dataset with a 4.6% improvement over the previous best PSNR, thus striking a better balance in reconstruction quality. The proposed model is lightweight and efficient, making it suitable for real-time applications in clinical and research settings. The code is available at https://github.com/hdy6438/Spec2VolCAMU-Net.
@@
journal: Biomedical Signal Processing and Control
@@ 1 Introduction
Recent advancements have leveraged deep learning to tackle this complex mapping task, requiring a comprehensive spatiotemporal feature extraction from EEG spectrograms and the synthesis of high-resolution fMRI volumes. Liu and Sajda pioneered the concept of ”neural transcoding” using coupled Convolutional Neural Networks (CNNs) to map between EEG and fMRI without explicit physiological priors [7], demonstrating feasibility on simulated data, yet their single-scale convolutions cannot capture discriminative cross-channel, cross-frequency patterns, leaving the model vulnerable to noise and acquisition differences, making its lacking robustness to noise and acquisition differences in real scenarios, and the lack of validation on real EEG-fMRI datasets raises concerns about its practical applicability. Later, Calhas et al. investigated autoencoders and Generative Adversarial Networks (GANs), later proposing models incorporating topographical attention mechanisms and Fourier features to explicitly model relationships between EEG electrodes, significantly improving synthesis performance [13]. However, the random Fourier encoding still entangles temporal and frequency cues, and its multi-branch architecture and random Fourier features lead to a huge hyperparameter space, requiring expensive Bayesian or grid search [14, 15]; More recently, transformer [16]-based architectures have been introduced, such as the Neural Transcoding Vision Transformer (NT-ViT) [17] by Lanzino et al., which utilizes a domain-matching module to align latent representations across modalities. However, this kind of Transformer-based model has a surge in computational complexity when the patch size is reduced or the number of layers is increased [18, 19], and the author also clearly pointed out that the output can only be used as an ”approximation of fMRI” rather than a substitute; Roos et al. proposed E2fNet [6], a U-Net-inspired CNN architecture demonstrating robust performance, particularly in structural similarity, but it still suffers from the limitations of CNNs, such as limited long-range dependencies and high computational costs. Besides, Roos et al. also proposed E2fGAN [6], a GAN-based model [20] that generates fMRI volumes from EEG spectrograms. However, GANs are known for their training instability and mode collapse [20, 21, 22], leading to suboptimal performance in some cases.
@@
Despite these advancements, existing EEG-to-fMRI synthesis methods face challenges in effective extracting discriminative time-frequency features that capture both local details and cross-channel/global patterns, and modeling long-range spatial dependencies in the fMRI volume without prohibitive memory or training-instability costs. To address the above challenges, we proposed Spec2VolCAMU-Net: First, we propose an EEG encoder named Multi-directional Time-Frequency Convolutional Attention Encoder (MD-TF-CAE) that captures both local and global features from EEG spectrograms. The MD-TF-CAE consists of three parallel convolutional paths: temporal convolution, frequency convolution, and joint time-frequency convolution. Then, the Vision-Mamba U-Net (VM-Unet) [23] decoder is used to replace self-attention with visual state space blocks (VSS Block) [24], reducing the complexity of sequence modeling to linear, significantly reducing video memory and inference time, and obtaining smoother gradients and enhanced convergence due to the lack of adversarial targets. With these targeted designs, Spec2VolCAMU-Net simultaneously improves accuracy, stability, and cross-dataset robustness without increasing the complexity of the architecture, providing an efficient and reliable new path for EEG-to-fMRI synthesis.
@@
Through comprehensive evaluation on multiple publicly available simultaneous EEG-fMRI datasets, including NODDI, Oddball, and CN-EPFL, Spec2VolCAMU-Net demonstrates superior performance compared to existing state-of-the-art approaches. The proposed Spec2VolCAMU-Net consistently achieves the highest Structural Similarity Index (SSIM) across all datasets, indicating exceptional structural fidelity in the reconstructed fMRI volumes. Specifically, Spec2VolCAMU-Net obtained SSIM values of $\textbf{0.693}\pm 0.048$ on NODDI, $\textbf{0.725}\pm 0.040$ on Oddball, and 0.788 on CN-EPFL, surpassing all compared models. Furthermore, Spec2VolCAMU-Net yielded the highest PSNR on the CN-EPFL dataset (23.837), highlighting its effectiveness in minimizing reconstruction errors. These quantitative metrics, complemented by qualitative visualizations, underscore Spec2VolCAMU-Net’s ability to generate highly realistic and accurate 3D fMRI volumes from EEG spectrogram inputs.
@@ 2.2.3 Vision-Mamba U-Net Decoder
In the fMRI image reconstruction stage, the aligned feature map $F$ is input into a Vision-Mamba U-Net (VM-UNet) [23] for multi-scale feature decoding and spatial reconstruction. The VM-UNet, as the decoder, integrates the local modeling capabilities of convolutional neural networks with the long-range modeling advantages of Mamba modules, demonstrating excellent image reconstruction and detail restoration capabilities. The decoder ultimately transforms the feature map with a channel dimension of $N$ into a three-dimensional voxel image $Y\in\mathbb{R}^{D\times H\times W}$, which is the synthesized fMRI output.
The VM-UNet decoder is a convolution-free architecture that replaces the traditional convolutional layers with Visual state space blocks (VSS Blocks) [24]. The VSS Block is a lightweight module that captures local features through 2D-Selective-Scan (SS2D) module, which is a variant of the Mamba module [23]. The SS2D module is designed to efficiently capture local features in the spatial domain while maintaining the long-range modeling capabilities of the Mamba module [25]. The SS2D module consists of three components: a scan expanding operation, an S6 block [25], and a scan merging operation. The scan expanding operation unfolds the input feature map along four different directions (top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right) into sequences. These sequences are then processed by the S6 block for feature extraction, ensuring that information from various directions is thoroughly scanned, thus capturing diverse features. Subsequently, the scan merging operation sums and merges the sequences from the four directions, restoring the output feature map to the same size as the input.
@@ 2.3 Loss Function
$\mathcal{L}_{total}=\lambda_{1}\cdot\mathcal{L}_{SSIM}+\lambda_{2}\cdot\mathcal{L}_{MSE}$ (1)
where $\lambda_{1},\lambda_{2}\in\mathbb{R}_{\geq 0}$ are weighting coefficients.
@@ 3.2 Dataset
The NODDI dataset is a resting-state EEG-fMRI dataset contains simultaneous recordings from 17 healthy adults who fixated on a white cross for about 11 min. EEG was acquired with a 64-channel 10-20 cap at 250 Hz, while whole-brain fMRI volumes were collected on a Siemens Avanto 1.5 T scanner ($TR=2.16s$, $30\times 64\times 64$ voxels, $3{\times}3{\times}$ 3 × 3 mm). Following common practice, two subjects with excessive motion were discarded, leaving 15 participants and about 4,110 paired EEG-fMRI samples after temporal alignment and windowing.
@@
The CN-EPFL consists of 20 subjects performing a speeded visual discrimination task with trial-by-trial confidence ratings. EEG was recorded at 5 kHz from 64 channels, and fMRI volumes were captured on a Siemens Trio 3 T (TR = 1.28 s, $54\times 108\times 108$ voxels, $2{\times}2{\times}$ 2 × 2 mm). To harmonise spatial resolution with the other benchmarks we follow earlier work and apply a DCT-based down-sampling to $30\times 64\times 64$, yielding 6 880 EEG-fMRI pairs.
@@ 3.3 Baseline Methods
    NT-ViT [17]. A Vision-Transformer generator that tokenises 3-D fMRI patches and EEG Mel-spectrogram patches, then enforces domain matching between their latent tokens. On NODDI and Oddball the authors report a 10×10$\times$10 × lower RMSE and 3.1×3.1$\times$3.1 × higher SSIM than earlier CNNs, setting a new state-of-the-art for transformer baselines.
@@ 3.5 Implementation Details
The Spec2VolCAMU-Net model is implemented using PyTorch and trained on a single NVIDIA GeForce RTX 4070 GPU. The model is trained for 50 epochs with a batch size of 16, using the AdamW optimizer with an initial learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-2}$. A cosine annealing schedule with hard restarts is applied, where the learning rate restarts every 10 epochs. The loss function is a combination of SSIM and MSE losses, with $\lambda_{1}=0.5$ and $\lambda_{2}=0.5$.
@@ 4 Results and Discussion
The proposed Spec2VolCAMU-Net model demonstrates a significant advancement in the field of EEG-to-fMRI reconstruction, consistently achieving state-of-the-art performance in structural similarity across multiple challenging public datasets. As detailed in Table 2, Spec2VolCAMU-Net markedly surpasses existing leading methods, including E2fNet, NT-ViT, and E2fGAN, in terms of the SSIM on the NODDI, Oddball, and CN-EPFL datasets. Specifically, the model achieved SSIM scores of $0.693\pm 0.048$ on NODDI, $0.725\pm 0.040$ on Oddball, and $0.788$ on CN-EPFL, representing improvements of 14.5%, 14.9%, and 16.9% respectively, over the previous best SSIM scores reported by E2fNet. This substantial enhancement in SSIM is critical as it indicates a superior capability to preserve the neuro-anatomical structure and spatial layout of the brain in the synthesized fMRI volumes, a key objective in cross-modal neuroimaging translation.
@@
To further illustrate the performance of Spec2VolCAMU-Net, we present qualitative visual comparisons of reconstructed fMRI volumes across all three datasets in Figures 3, 4, and 5. The first column shows the ground truth (GT) fMRI volume, while the subsequent columns display the results from E2fNet, E2fGAN, and Spec2VolCAMU-Net.
Figure 3: Visual comparison of reconstructed fMRI volumes from Spec2VolCAMU-Net and certain baseline methods on the CN-EPFL dataset. The first column shows the ground truth (GT) fMRI volume, while the subsequent columns display the results from E2fNet, E2fGAN, and Spec2VolCAMU-Net.

### v2

@@ Abstract
High-resolution functional magnetic resonance imaging (fMRI) is essential for mapping human brain activity; however, it remains costly and logistically challenging. If comparable volumes could be generated directly from widely available scalp electroencephalography (EEG), advanced neuroimaging would become significantly more accessible. Existing EEG-to-fMRI generators rely on plain Convolutional Neural Networks (CNNs) that fail to capture cross-channel time-frequency cues or on heavy transformer/Generative Adversarial Network (GAN) decoders that strain memory and stability. To address these limitations, we propose Spec2VolCAMU-Net, a lightweight architecture featuring a Multi-directional Time-Frequency Convolutional Attention Encoder for rich feature extraction and a Vision-Mamba U-Net decoder that uses linear-time state-space blocks for efficient long-range spatial modelling. We frame the goal of this work as establishing a new state of the art in the spatial fidelity of single-volume reconstruction, a foundational prerequisite for the ultimate aim of generating temporally coherent fMRI time series. Trained end-to-end with a hybrid SSI-MSE loss, Spec2VolCAMU-Net achieves state-of-the-art fidelity on three public benchmarks, recording Structural Similarity Index (SSIM) of 0.693 on NODDI, 0.725 on Oddball and 0.788 on CN-EPFL, representing improvements of 14.5%, 14.9%, and 16.9% respectively over previous best SSIM scores. Furthermore, it achieves competitive Signal-to-Noise Ratio (PSNR) scores, particularly excelling on the CN-EPFL dataset with a 4.6% improvement over the previous best PSNR, thus striking a better balance in reconstruction quality. The proposed model is lightweight and efficient, making it suitable for real-time applications in clinical and research settings. The code is available at https://github.com/hdy6438/Spec2VolCAMU-Net.
@@
journal: Journal of Neural Engineering
@@ 1 Introduction
Early pioneering work in this field utilized statistical machine learning methods. This foundational research often fell into two main paradigms: EEG-informed fMRI analysis [13] and joint independent component analysis (ICA) [14]. The former is a hypothesis-driven approach where features derived from EEG, such as event-related potentials or epileptic spikes, are used as regressors in a General Linear Model (GLM) to analyze fMRI data [15]. This allowed researchers to pinpoint fMRI BOLD responses that were spatiotemporally correlated with EEG events. In contrast, joint ICA is a data-driven technique that decomposes both EEG and fMRI data to identify shared, spatially independent components [16], which helps uncover common brain networks and non-obvious relationships between the two modalities, especially in resting-state studies. These foundational methods paved the way for more targeted predictive models. For example, Meir-Hasson et al. proposed a model based on a time/frequency representation with varying time delays to predict amygdala activity from a single EEG electrode [17]. This work demonstrated the feasibility of extracting meaningful information from limited EEG data to infer activity in deep brain structures. Similarly, Cury et al. introduced a sparse regression model to predict fMRI-derived neurofeedback scores from EEG signals alone, demonstrating the potential of statistical approaches to bridge the modalities [13]. The use of a sparsity constraint in their model allowed for the selection of the most relevant EEG features, enhancing model interpretability and preventing overfitting. These methods laid the groundwork for leveraging the temporal characteristics of EEG to inform fMRI signals, providing crucial theoretical and methodological precedents for subsequent deep learning-based approaches.
@@
Recent advancements have leveraged deep learning to tackle this complex mapping task, requiring a comprehensive spatiotemporal feature extraction from EEG spectrograms and the synthesis of high-resolution fMRI volumes. Liu and Sajda pioneered the concept of ”neural transcoding” using coupled Convolutional Neural Networks (CNNs) to map between EEG and fMRI without explicit physiological priors [7], demonstrating feasibility on simulated data, yet their single-scale convolutions cannot capture discriminative cross-channel, cross-frequency patterns, leaving the model vulnerable to noise and acquisition differences, making its lacking robustness to noise and acquisition differences in real scenarios, and the lack of validation on real EEG-fMRI datasets raises concerns about its practical applicability. Later, Calhas et al. investigated autoencoders and Generative Adversarial Networks (GANs), later proposing models incorporating topographical attention mechanisms and Fourier features to explicitly model relationships between EEG electrodes, significantly improving synthesis performance [18]. However, the random Fourier encoding still entangles temporal and frequency cues, and its multi-branch architecture and random Fourier features lead to a huge hyperparameter space, requiring expensive Bayesian or grid search [19, 20]; More recently, transformer [21]-based architectures have been introduced, such as the Neural Transcoding Vision Transformer (NT-ViT) [22] by Lanzino et al., which utilizes a domain-matching module to align latent representations across modalities. However, this kind of Transformer-based model has a surge in computational complexity when the patch size is reduced or the number of layers is increased [23, 24], and the author also clearly pointed out that the output can only be used as an ”approximation of fMRI” rather than a substitute; Roos et al. proposed E2fNet [6], a U-Net-inspired CNN architecture demonstrating robust performance, particularly in structural similarity, but it still suffers from the limitations of CNNs, such as limited long-range dependencies and high computational costs. Besides, Roos et al. also proposed E2fGAN [6], a GAN-based model [25] that generates fMRI volumes from EEG spectrograms. However, GANs are known for their training instability and mode collapse [25, 26, 27], leading to suboptimal performance in some cases. More recently, some deep learning approaches have explicitly incorporated the time-series nature of the data. For instance, NeuroBOLT, a multi-dimensional transformer-based framework, translates raw EEG time series to fMRI time series by learning representations across temporal, spatial, and spectral domains, enabling more generalized models [28]. Another work by Li et al. proposed a novel architecture using a Sinusoidal Representation Network (SIREN) to predict fMRI signals directly from multi-channel EEG without explicit feature engineering, focusing on learning frequency information implicitly [29].
@@
Despite these advancements, existing EEG-to-fMRI synthesis methods face challenges in effective extracting discriminative time-frequency features that capture both local details and cross-channel/global patterns, and modeling long-range spatial dependencies in the fMRI volume without prohibitive memory or training-instability costs. To address the above challenges, we proposed Spec2VolCAMU-Net: First, we propose an EEG encoder named Multi-directional Time-Frequency Convolutional Attention Encoder (MD-TF-CAE) that captures both local and global features from EEG spectrograms. The MD-TF-CAE consists of three parallel convolutional paths: temporal convolution, frequency convolution, and joint time-frequency convolution. Then, the Vision-Mamba U-Net (VM-Unet) [30] decoder is used to replace self-attention with visual state space blocks (VSS Block) [31], reducing the complexity of sequence modeling to linear, significantly reducing video memory and inference time, and obtaining smoother gradients and enhanced convergence due to the lack of adversarial targets. With these targeted designs, Spec2VolCAMU-Net simultaneously improves accuracy, stability, and cross-dataset robustness without increasing the complexity of the architecture, providing an efficient and reliable new path for EEG-to-fMRI synthesis, achieving a new state of the art in the spatial fidelity of single-volume reconstruction, which is a foundational prerequisite for the ultimate aim of generating temporally coherent fMRI time series.
@@
Through comprehensive evaluation on multiple publicly available simultaneous EEG-fMRI datasets, including NODDI, Oddball, and CN-EPFL, Spec2VolCAMU-Net demonstrates superior performance compared to existing state-of-the-art approaches. The proposed Spec2VolCAMU-Net consistently achieves the highest Structural Similarity Index (SSIM) across all datasets, indicating exceptional structural fidelity in the reconstructed fMRI volumes. Specifically, Spec2VolCAMU-Net obtained SSIM values of $\textbf{0.693}\pm 0.048$ on NODDI, $\textbf{0.725}\pm 0.040$ on Oddball, and 0.788 on CN-EPFL, surpassing all compared models. Furthermore, Spec2VolCAMU-Net yielded the highest Signal-to-Noise Ratio (PSNR) on the CN-EPFL dataset (23.837), highlighting its effectiveness in minimizing reconstruction errors. These quantitative metrics, complemented by qualitative visualizations, underscore Spec2VolCAMU-Net’s ability to generate highly realistic and accurate 3D fMRI volumes from EEG spectrogram inputs.
@@ 2.2.3 Vision-Mamba U-Net Decoder
For the final synthesis of the fMRI volume, the aligned feature map $F$ is processed by a Vision-Mamba U-Net (VM-UNet) [30] decoder. This decoder is responsible for multi-scale feature decoding and spatial reconstruction, ultimately transforming the $N$-channel feature map into the final three-dimensional fMRI volume $Y\in\mathbb{R}^{D\times H\times W}$. The architecture of the VM-UNet is designed to synergize the capacity for local feature modeling, typically associated with CNNs, with the efficient long-range dependency modeling of Mamba modules, making it highly effective for detailed image reconstruction.
Instead of relying on traditional convolutions, the VM-UNet decoder is built upon Visual State Space (VSS) blocks [31]. At the core of each VSS block is the 2D-Selective-Scan (SS2D) module, a Mamba variant adapted for vision tasks [30, 32]. The SS2D mechanism is engineered to capture spatial information efficiently. It operates by first expanding the input feature map into four separate sequences by scanning it in four different directions (e.g., top-left to bottom-right). Each sequence is then processed by a structured state-space model (S6 block) [32] to extract features. Finally, a merging step combines the information from all four directional scans, producing an output feature map of the same dimensions as the input but enriched with multi-directional context.
@@ 2.3 Loss Function
$\mathcal{L}_{total}=\lambda_{SSIM}\cdot\mathcal{L}_{SSIM}+\lambda_{MSE}\cdot\mathcal{L}_{MSE}$ (1)
where $\lambda_{SSIM},\lambda_{MSE}\in\mathbb{R}_{\geq 0}$ are weighting coefficients, and $\lambda_{SSIM}+\lambda_{MSE}=1$.
@@ 3.2 Dataset
The NODDI dataset is a resting-state EEG-fMRI dataset contains simultaneous recordings from 17 healthy adults who fixated on a white cross for about 11 min. EEG was acquired with a 64-channel 10-20 cap at 250 Hz, while whole-brain fMRI volumes were collected on a Siemens Avanto 1.5 T scanner ($TR=2.16s$, $30\times 64\times 64$ voxels, $3{\times}3{\times}3$ mm). Following common practice, two subjects with excessive motion were discarded, leaving 15 participants and about 4,110 paired EEG-fMRI samples after temporal alignment and windowing.
@@
The CN-EPFL consists of 20 subjects performing a speeded visual discrimination task with trial-by-trial confidence ratings. EEG was recorded at 5 kHz from 64 channels, and fMRI volumes were captured on a Siemens Trio 3 T (TR = 1.28 s, $54\times 108\times 108$ voxels, $2{\times}2{\times}2$ mm). To harmonise spatial resolution with the other benchmarks we follow earlier work and apply a DCT-based down-sampling to $30\times 64\times 64$, yielding 6 880 EEG-fMRI pairs.
@@ 3.3 Baseline Methods
    NT-ViT [22]. A Vision-Transformer generator that tokenises 3-D fMRI patches and EEG Mel-spectrogram patches, then enforces domain matching between their latent tokens. On NODDI and Oddball the authors report a $10\times$ lower RMSE and $3.1\times$ higher SSIM than earlier CNNs, setting a new state-of-the-art for transformer baselines.
@@ 3.5 Implementation Details
The Spec2VolCAMU-Net model is implemented using PyTorch and trained on a single NVIDIA GeForce RTX 4070 GPU. The model is trained for 50 epochs with a batch size of 16, using the AdamW optimizer with an initial learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-2}$. A cosine annealing schedule with hard restarts is applied, where the learning rate restarts every 10 epochs. The loss function is a combination of SSIM and MSE losses, weighted by $\lambda_{SSIM}$ and $\lambda_{MSE}$.
Figure 2: Trade-off between SSIM and PSNR on the validation set for different values of $\lambda_{SSIM}$ and $\lambda_{MSE}$. The optimal balance is achieved at $\lambda_{SSIM}=0.5$ and $\lambda_{MSE}=0.5$.
To determine the optimal values of $\lambda_{SSIM}$ and $\lambda_{MSE}$, a grid search was performed over the range $[0,1]$ with a step size of 0.2, ensuring that $\lambda_{SSIM}+\lambda_{MSE}=1$. The model was evaluated on a CN-EPFL dataset using SSIM and PSNR as metrics, and the combination of $\lambda_{SSIM}=0.5$ and $\lambda_{MSE}=0.5$ yielded the best overall performance across all datasets. The trade-off between SSIM and PSNR for different values of $\lambda_{SSIM}$ and $\lambda_{MSE}$ is illustrated in Figure 2.
@@
3.6 Ablation Study
To validate the effectiveness of the proposed components in Spec2VolCAMU-Net, we conduct an ablation study on the CN-EPFL dataset. We systematically remove or replace key components of the model and evaluate the impact on performance using SSIM and PSNR metrics. Specifically, we replace the Vision-Mamba U-Net decoder with a standard U-Net architecture, and remove each of the three convolutional paths in the MD-TF-CAE (temporal, frequency, and joint time-frequency), as well as the multi-head self-attention module.
@@ 4.1 Performance Comparison with State-of-the-Art Methods
The results for other methods are taken from their respective papers.
The proposed Spec2VolCAMU-Net model demonstrates a significant advancement in the field of EEG-to-fMRI reconstruction, consistently achieving state-of-the-art performance in structural similarity across multiple challenging public datasets. As detailed in Table 2, Spec2VolCAMU-Net markedly surpasses existing leading methods, including E2fNet, NT-ViT, and E2fGAN, in terms of the SSIM on the NODDI, Oddball, and CN-EPFL datasets. Specifically, the model achieved SSIM scores of $0.693\pm 0.048$ on NODDI, $0.725\pm 0.040$ on Oddball, and $0.788$ on CN-EPFL, representing improvements of 14.5% ($p\ll 0.001$, 95% CI of difference [0.053,0.122]), 14.9% ($p\ll 0.001$, 95% CI of difference [0.065,0.123]), and 16.9% respectively, over the previous best SSIM scores reported by E2fNet. This substantial enhancement in SSIM is critical as it indicates a superior capability to preserve the neuro-anatomical structure and spatial layout of the brain in the synthesized fMRI volumes, a key objective in cross-modal neuroimaging translation.
@@
4.2 Qualitative Visual Comparisons
To further illustrate the performance of Spec2VolCAMU-Net, we present qualitative visual comparisons of reconstructed fMRI volumes across all three datasets in Figures 4, 5, and 6. The first column shows the ground truth (GT) fMRI volume, while the subsequent columns display the results from E2fNet, E2fGAN, and Spec2VolCAMU-Net.
Figure 4: Visual comparison of reconstructed fMRI volumes from Spec2VolCAMU-Net and certain baseline methods on the CN-EPFL dataset. The first column shows the ground truth (GT) fMRI volume, while the subsequent columns display the results from E2fNet, E2fGAN, and Spec2VolCAMU-Net. The brain regions that improve the most with Spec2VolCAMU-Net are highlighted with red boxes.